# Nonlinear Permuted Granger Causality

### v1

@@ 5 Application Study
The out-of-sample permutation framework is applied to neuronal responses to acoustic stimuli in the primary auditory cortex of an anesthetized (ketamine–medetomidine) rat. The data, taken from the Collaborative Research in Computational Neuroscience data sharing website, consists of in vivo whole-cell recordings (mV) sampled at 4kHz in response to natural sound fragments (Asari et al., 2009; Machens et al., 2004). We examine experimental trial 60 from the 050802mw03 data and isolate a time region of interest between 11 and 15 seconds (partially shown in Figures 2D and 2E of Machens et al. (2004)). The subset of data corresponds to the recordings in response to a jaguar (Panthera onca) mating call sound fragment presented at 97.656 kHz; we resample to a frequency of 4kHz, matching that of the response. To construct a simplistic non-causal scenario, we include additional sound fragments of a Humpback whale (Megaptera novaeangliae) and Knudsen’s frog (Leptodactylus knudseni) that are played in other trials throughout the experiment. Figure 3 displays the whole-cell recordings to the jaguar mating call and all sound fragments examined. We refer to Machens et al. (2004) for additional detail on data collection and experimental methods.
@@
Figure 3: In vivo whole-cell recordings (mV) from the contralateral (left) primary auditory cortex of an anesthetized rat (top) in response to a jaguar mating call acoustic stimulus fragment resampled at 4kHz (bottom). Humpback whale and Knudsen’s frog acoustic stimuli are included as non-causal examples for the methodology.
@@
We examine the performance of the Granger causal detection methodology by formulating a farcical example in which any suitable method should be able to recover the correct causal structure. The methods are expected to flag the jaguar sound fragment as causal, while correctly labelling the other two included fragments non-causal. We isolate lagged information from 30ms to 40ms, consistent with the measured the half-maximal synaptic conductance in Wehr and Zador (2003). One acoustic stimulus fragment is chosen as the covariate set $\mathbf{X}$ and the others are additional variables $\mathbf{Z}$ in the model. Lagged values of the response are included up to 10ms. We implement the NPGC method with $M=400$ permutations, $K=5$ cross validation folds, $\mathscr{R}=50$ random initializations, and due to the large number of variables, we increase the feature dimension to $N=250$. The comparator penalized optimization methods are examined at the same dimension and a variety of penalty values. We note the challenge in selection of the penalty $\lambda$ and emphasize that this value must be chosen prior to analysis and not fine tuned after the fact to produce a desired (already known) result. Results for the fabricated, naive scenario are compiled in Figure 4, and we demonstrate that the permutation based methodology extracts the correct relationship between the jaguar mating call sound fragment and the whole-cell recording response where the penalized optimization approaches struggle.
@@
Figure 4: Estimated Granger causal relationship between each acoustic stimulus and primary auditory cortex response to the jaguar mating call sound fragment (NPGC quantiles $\hat{Q}_{M}=0.0025,1.000,0.4425,$ from top to bottom). The NPGC method recovers the correct causal structure, while the lasso-based cMLP methods fail to isolate the jaguar stimulus.
6 Conclusion
@@
The NPGC methodology illustrates the use of the permutation-based shift from in-sample to out-of-sample testing. Permutation tests are used widely in literature, and Nauta et al. (2019) implements a decision rule similar to NPGC. We explicitly define Granger causality in this framework, and provide a theoretical analysis of the corresponding variance estimates and decision rule. This shift allows for control in identifying useful functional relationships when overfitting from an artificial network becomes a concern, and removes the burden of specifying regularization parameters that have a major impact on the observed outcome. Once a Granger causal outcome has been determined, penalized optimization techniques provide a practical screening method for identifying the prominent relationships, but they do not effectively estimate the global presence or absence of functional connectivity.
@@
We show the ability of the permutation framework to detect the presence of functional connectivity, and to provide a safeguard against misidentification of data-specific noise as functional dependence when the artificial network is overfit. Alternative formulations of the featurizing function $\Psi$ in Equation 4 can utilize the same methodological structure and retain the theoretical guarantees provided the transformation meets the required conditions listed at the beginning of Section 3 and the construction allows for row-wise permutation without breaking the dependence structure across variables in $\mathbf{X}$. Misspecification of the dimension of the feature space should only produce additional false negative results as long as $N$ is less than the number of data points used in model estimation. We suggest using a large enough dimension to capture any potential nonlinear dependence structure that may exist in the dataset. We also note that the NPGC method exhibits slight undercoverage for a chosen level of test, but argue that minor deviations in this realm are not of great concern if the result is correctly interpreted as a potential functional connection and not an outright causal effect. The permutation method circumvents the need for a penalty parameter selection, and provides ease of extension to multiple testing problems and control of family-wise error rate.
@@
We caution against potential misuses of Granger causality including, but not limited to, repeated application to subsets of a selected covariate group without adequate Type 1 error correction, disregard for the conditional nature of the inferential conclusion, attempts at individual rather than collective covariate inference (unless specification of the model is complete and exact), inclusion of a covariate without careful scientific or logical reasoning, and its use as an outright mechanism for identifying causal relationships rather than predictive links for future study. We also recommend prudent selection of the length of lag response included in the covariate matrix and recognize that two values may produce different results. In nonlinear processes, selection of the relevant history of the response for inclusion in the model remains an area of future study.
@@
Nonlinear Permuted Granger Causality
@@
: Tables and figures of simulation and application results not presented in the main paper. (PDF)
Algorithms & Pseudocode:
Presentation of algorithms mentioned in Section 2 in the form of pseudocode. See main text for explanation. (PDF)
Theory Supplement:
Conditions for theoretical results and proofs for theorems presented in the paper. (PDF)
Simulation data:
Simulated data used to assess performance of methods available at
https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.RData, .npy, and .csv files)
Simulation code:
Code used to generate results and figures discussed in paper available at https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.R, .cpp, and.py files)
Application data:
Data used for applications Section 5 available at
https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.mat file)
Application code:
Code used to assess the methods in the application study available at https://github.com/noahgade/NonlinearPermutedGrangerCausality. (.R, .py, and .m files)
7 Additional Tables & Figures
@@
(a) TAR(2) response with zero Granger causal variables included in additional set $\mathbf{Z}$.
@@
(b) TAR(2) response with two Granger causal variables included in additional set $\mathbf{Z}$.
Table 5: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for TAR(2) simulations.
@@
(a) Lorenz-96 response with zero Granger causal variables included in additional set $\mathbf{Z}$.
@@
(b) Lorenz-96 response with two Granger causal variables included in additional set $\mathbf{Z}$.
Table 6: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for Lorenz-96 simulations.
Figure 5: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations with zero (top) and two (bottom) Granger causal variables included in additional set $\mathbf{Z}$.
Figure 6: Type 1 Error control by reference quantile $\alpha$ for Lorenz-96 simulations with zero (top) and two (bottom) Granger causal variables included in additional set $\mathbf{Z}$.
Figure 7: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations without Lorenz-96 group included to demonstrate adherence to theoretical results. Zero Granger causal variables (left) and two Granger causal variables (right) included in the additional set $\mathbf{Z}$. All data sets are $T=1000$.
8 Algorithms for Methodology
@@
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # permutations $M$; # random initializations $\mathscr{R}$; feature space dimension $N$; # cross validation folds $K$
@@
6:     $\mathbf{H}_{\omega,m,r}\leftarrow\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{m,\omega}\right]\mathbf{W}_{r}\right)$
@@
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # random initializations $\mathscr{R}$; # cross validation folds $K$
@@
9 Theory Supplement
We begin with the additional conditions for theoretical results presented in Section 3.
9.1 Conditions for Theoretical Results
Define the functional transformation $\Psi$ in Equation 4 with element-wise nonlinear activation function $g=\tanh$. We derive the results for a generic $g$ under the constraints of Condition 4.
Condition 4.
The nonlinear function activation function $g$ in Equation 4 is bounded, $g:\mathbb{R}\rightarrow\left[a,b\right]$ for some $a<b$, such that for any $x\in\mathbb{R}$, $|g(x)|\leq G=\max\left\{|a|,|b|\right\}$ and $G<\infty$.
Condition 5.
Let $\mathcal{W}$ be the space of all element-wise random initializations, $w_{ij}\sim\mathcal{N}(0,1)$, such that for all $\mathbf{W}_{r}\in\mathcal{W}$, the matrix $\mathbf{W}_{r}$ is full rank and $\Psi$ generates a feature matrix that is full rank with a finite condition number. For permutation $m$, realization $\omega$, and random initialization $r$,
@@ 9.1 Conditions for Theoretical Results
For the matrix $\mathbf{H}_{m,\omega,r}=(h_{m,\omega,r,ij})$, define $G$ as the maximal element from Condition 4 and the average squared entry as $\bar{h}_{m,\omega,r}^{2}$.
Condition 6.
The following expected values exist and are finite:
@@
Combining Conditions 4, 5, and 6, $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$. We define $f$ as the true functional relationship between the response and the unpermuted predictor matrix for all $\omega\in\Omega$, $\mathbf{Y}_{\omega}=f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)+\bm{\varepsilon}_{\omega}$, and the approximating model form for permutations $m=1,\ldots,M$ and initializations $r=1,\ldots,\mathscr{R}$.
$\mathbf{Y}_{\omega}=\mathbf{H}_{m,\omega,r}\mathbf{B}_{m,\omega,r}+\mathbf{U}_{m,\omega,r}$ (30)
Condition 7.
For every $\eta>0$, there exists a fixed $N$, where $1+\gamma d+q+p\leq N<\infty$, such that as the number of random initializations $\mathscr{R}\rightarrow\infty$,
$\sup_{\omega\in\Omega}\left\|\frac{1}{\mathscr{R}}\sum_{r=1}^{\mathscr{R}}\mathbf{H}_{1,\omega,r}\mathbf{B}_{1,\omega,r}-f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)\right\|<\eta,$ (31)
where $\mathbf{B}_{1,\omega,r}$ is the true coefficient matrix of feature space $r$ for the unpermuted covariate set, $\mathbf{H}_{1,\omega,r}=\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\mathbf{W}_{r}\right)$.
We note that Condition 7 specifically pertains to the $f$ piece of the true functional relationship; we make no assumption of the closeness of a transformed response $\mathbf{H}_{1,\omega,r}\mathbf{B}_{1,\omega,r}$ and the true values $\mathbf{Y}_{\omega}$ if the predictors themselves are not reliable and the entries of $\bm{\varepsilon}_{\omega}$ are large.
9.2 Proofs
Lemma 1.
Under the conditions listed in Section 3 and the supplement, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$.
The proof of Lemma 1 begins with the estimate for one test set with a single individual random initialization given the variation measure $\mathbf{S}_{m,\omega,r}$. We aggregate these expectations over the sets $k=1,\ldots,K$, realizations $r=1,\ldots,\mathscr{R}$ and the $\varphi\geq 1$ observed datasets in $\Omega_{\text{obs}}$.
Proof of Lemma 1.
@@ 9.2 Proofs
We examine the matrices $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ under Conditions 4, 5, 6, and 7. Recall $T_{k}$ is the number of rows in the test set, and $T_{-k}$ the number in the training set. We obtain a fixed and finite $N$ from Condition 7, and $\text{rank}(\mathbf{H}_{m,\omega,r,-k})=N$ under Equation 25 from Condition 5. We label the singular values of the matrices $\sigma_{i}(\mathbf{H}_{m,\omega,r,-k})>0$ and $\sigma_{i}(\mathbf{H}_{m,\omega,r,k})\geq 0$ for $i=1,\ldots,N$. For any matrix $\mathbf{P}$, $\left\|\mathbf{P}\right\|_{F}=\sqrt{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|p_{ij}|^{2}}=\sqrt{\sum_{i=1}^{\text{rank}(\mathbf{P})}\sigma_{i}(\mathbf{P})^{2}}$. From Condition 4, each entry in $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ lies in a bounded interval, $h_{m,\omega,r,-k,ij},h_{m,\omega,r,k,ij}\in[a,b]$ such that $|h_{m,\omega,r,-k,ij}|,|h_{m,\omega,r,k,ij}|\leq G$, and we use this to establish the upper bounds for the maximal singular values shown in Equations 47 and 48. To write the minimum values, we define the average squared matrix entry as in Condition 6, and note the combination of Conditions 4, 5, and 6 yields $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$.
@@
Under Condition 5, $\mathbf{H}_{m,\omega,r,-k}$ has a finite condition number $\kappa\left(\mathbf{H}_{m,\omega,r,-k}\right)\leq\kappa_{\max}$, and we can use the bounds in Equations 47 and 48 to obtain bounds for the minimum singular values. The rank of $\mathbf{H}_{m,\omega,r,k}$ is not necessarily $N$, as $T_{k}$ can be less than $N$, but it is at least $\min\left\{T_{k},N\right\}$.
@@
We return to the expectation equation given in Equation 32 with the two bounds from Equations 63 and 72.
@@
Thus, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$. ∎
Lemma 2.
Under the conditions listed in Section 3 and the supplement, $\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)=\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$.
For use in the proof of Lemma 2, we first state and prove Remark 1, and then proceed like in the proof of Lemma 1.
Remark 1.
@@
with the last line equal to $(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$. Thus, $\text{tr}^{2}\left(\mathbf{P}\right)\leq(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$. ∎
Proof of Lemma 2.
We isolate the quantity $\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)$ like in the proof of Lemma 1 and examine the conditional variance $\text{Var}\left(\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$. The unconditional variance follows from the law of total variance and from the definition in Equation 6.

### v2

@@ 5 Application Study
The out-of-sample permutation framework is applied to neuronal responses to acoustic stimuli in the primary auditory cortex of an anesthetized (ketamine–medetomidine) rat. The data, taken from the Collaborative Research in Computational Neuroscience data sharing website, consists of in vivo whole-cell recordings (mV) sampled at 4kHz in response to natural sound fragments (Machens et al., 2004; Asari et al., 2009). A time region of interest between 11 and 15 seconds is isolated from experimental trial 60 of the 050802mw03 data (partially shown in Figures 2D and 2E of Machens et al. (2004)). The subset of data corresponds to the recordings in response to a jaguar (Panthera onca) mating call sound fragment presented at 97.656 kHz; this is resampled to a frequency of 4kHz matching that of the response. Additional sound fragments of a Humpback whale (Megaptera novaeangliae) and Knudsen’s frog (Leptodactylus knudseni), played in other trials throughout the experiment, are included to construct a simplistic, non-causal scenario. Figure 3 displays the whole-cell recordings to the jaguar mating call and all sound fragments examined. Machens et al. (2004) contains additional detail on data collection and experimental methods.
@@
Figure 3: Top: In vivo whole-cell recordings (mV) from the contralateral (left) primary auditory cortex of an anesthetized rat. Bottom: Jaguar mating call acoustic stimulus fragment resampled at 4kHz. Humpback whale and Knudsen’s frog acoustic stimuli are included as non-causal examples for the methodology.
@@
Performance of the Granger causal detection methodology is evaluated by formulating a farcical example in which any suitable method should be able to recover the correct causal structure. The methods are expected to flag the jaguar sound fragment as causal, while correctly labelling the other two included fragments non-causal. Lagged information from 30ms to 40ms is isolated for inclusion, consistent with the measured the half-maximal synaptic conductance in Wehr and Zador (2003). One acoustic stimulus fragment is chosen as the covariate set $\mathbf{X}$ and the others are additional variables $\mathbf{Z}$ in the model. Lagged values of the response are included up to 10ms. The NPGC method is implemented with $M=400$ permutations, $K=5$ cross validation folds, $\mathscr{R}=50$ randomly generated FNNs, and a feature dimension of $N=250$ due to the large number of covariates. The comparator penalized optimization methods are examined at the same dimension and a variety of penalty values. The selection of the penalty $\lambda$ presents a major challenge, and this value must be chosen prior to analysis; it should not be fine-tuned after the fact to produce a desired (already known) result. Results for the fabricated, naive scenario are compiled in Figure 4.
@@
Figure 4: Estimated Granger causal relationship between each acoustic stimulus and primary auditory cortex response. NPGC quantiles $\hat{Q}_{M}=0.0025,1.000,0.4425,$ from top to bottom. The NPGC method recovers the correct causal structure, while the lasso-based cMLP methods fail to isolate the jaguar stimulus.
Permutation based methodology extracts the correct relationship between the jaguar mating call sound fragment and the whole-cell recording response, and the penalized variable selection approaches do not.
6 Discussion
@@
The NPGC methodology illustrates the use of the permutation-based shift from in-sample to out-of-sample testing. Permutation tests are used widely in literature, and Nauta et al. (2019) implement a decision rule similar to NPGC. This manuscript explicitly defines Granger causality in this framework, and it provides a theoretical analysis of the corresponding variance estimates and decision rule. The shift allows for control in identifying useful functional relationships when overfitting from an artificial network becomes a concern, and removes the burden of specifying regularization parameters that have a major impact on the observed outcome. Once a Granger causal outcome has been determined, penalized variable selection techniques provide a practical screening method for identifying the prominent relationships, but they do not effectively estimate the global presence or absence of functional connectivity.
@@
The permutation framework is able to detect the presence of functional connectivity, and provide a safeguard against misidentification of data-specific noise as functional dependence when the artificial network is overfit. Alternative formulations of the featurizing function $\Psi$ in Equation 10 can utilize the same methodological structure and retain the theoretical guarantees provided the transformation meets the listed conditions and the construction allows for row-wise permutation without breaking the dependence structure across variables in $\mathbf{X}$. Misspecification of the dimension of the feature space should only produce additional false negative results as long as $N$ is less than the number of data points used in model estimation. Using a large enough dimension to capture any potential nonlinear dependence structure that may exist in the dataset is suggested. The NPGC method may exhibit slight undercoverage for a chosen level of test, but minor deviations in this realm are not of great concern if the result is correctly interpreted as a potential functional connection and not an outright causal effect. The permutation method circumvents the need for a penalty parameter selection, and provides ease of extension to multiple testing problems and control of family-wise error rate.
@@
Potential misuses of Granger causality include, but are not limited to, repeated application to subsets of a selected covariate group without adequate Type 1 error correction, disregard for the conditional nature of the inferential conclusion, attempts at individual rather than collective covariate inference (unless model specification is complete and exact), inclusion of a covariate without careful scientific or logical reasoning, and use as an outright mechanism for identifying causal relationships rather than predictive links for future study. Prudent selection of the length of lag response included in the covariate matrix is required, and two values may produce different results. In nonlinear processes, selection of the relevant history of the response for inclusion in the model remains an area of future study.
@@
Title
Additional Tables & Figures:
Tables and figures of simulation and application results not presented in the main text.
Algorithms & Pseudocode:
Algorithms and pseudocode for methodology presented in the main text.
Theory Supplement:
Additional conditions for theoretical results, finite sample behavior, and proofs to theorems presented in the main text.
Code & Supporting Material:
Files (.RData, .npy, and .csv) used to assess performance of change point methods, code (.R, .cpp, and .py) used to generate results and figures, and data (.mat) and code (.R, .py, and .m) used to generate output in Section 5 can be found at
[…]
@@
Tables
@@
Table 6: TAR(2) simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for zero Granger causal variables included in additional set $\mathbf{Z}$.
@@
Table 7: TAR(2) simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for two Granger causal variables included in additional set $\mathbf{Z}$.
@@
Table 8: Lorenz-96 simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for zero Granger causal variables included in additional set $\mathbf{Z}$.
@@
Table 9: Lorenz-96 simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2) for two Granger causal variables included in additional set $\mathbf{Z}$.
Figures
Figure 5: Type 1 error control for TAR(2) simulations by reference quantile $\alpha$ with zero Granger causal variables included in additional set $\mathbf{Z}$.
Figure 6: Type 1 error control for TAR(2) simulations by reference quantile $\alpha$ with two Granger causal variables included in additional set $\mathbf{Z}$.
Figure 7: Type 1 error control for Lorenz-96 simulations by reference quantile $\alpha$ with zero Granger causal variables included in additional set $\mathbf{Z}$.
Figure 8: Type 1 error control for Lorenz-96 simulations by reference quantile $\alpha$ with two Granger causal variables included in additional set $\mathbf{Z}$.
Type 1 error control for two-group TAR(2) simulations by reference quantile $\alpha$ simulations without Lorenz-96 group included. Left: Zero Granger causal variables. Right: Two Granger causal variables included in the additional set $\mathbf{Z}$. All data sets are $T=1000$.
Algorithms & Pseudocode
@@
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # permutations $M$; # random featurizations $\mathscr{R}$; feature space dimension $N$; # cross-validation folds $K$
@@
6:     $\mathbf{H}_{m,\omega,r}\leftarrow\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{m,\omega}\right]\mathbf{W}_{r}\right)$
@@
Inputs: $\left(\mathbf{X},\mathbf{Y},\mathbf{Z}\right)_{\omega}$ for all $\varphi$ realizations $\omega\in\Omega_{\text{obs}}$; lag selection $\gamma$; # random FNN generations $\mathscr{R}$; # cross-validation folds $K$
@@
Theory Supplement
Additional Conditions for Theoretical Results
Theoretical results are derived for a generic activation function $g$ in Equation 10 with the constraints of Condition Additional Conditions for Theoretical Results. {condition} The nonlinear function activation function $g$ in Equation 10 is bounded, $g:\mathbb{R}\rightarrow\left[a,b\right]$ for some $a<b$, such that for any $x\in\mathbb{R}$, $|g(x)|\leq G=\max\left\{|a|,|b|\right\}$ and $G<\infty$. {condition} Let $\mathcal{W}$ be the space of all element-wise randomly generated FNNs, $w_{ij}\sim\mathcal{N}(0,1)$, such that for all $\mathbf{W}_{r}\in\mathcal{W}$, the matrix $\mathbf{W}_{r}$ is full rank and generates a feature matrix that is full rank with a finite condition number. For permutation $m$, realization $\omega$, and random featurization $r$,
@@
For the matrix $\mathbf{H}_{m,\omega,r}=(h_{m,\omega,r,ij})$, define $G$ as the maximal element from Condition Additional Conditions for Theoretical Results and the average squared entry as $\bar{h}_{m,\omega,r}^{2}$. {condition} The following expected values exist and are finite:
@@
Combining Conditions Additional Conditions for Theoretical Results, Additional Conditions for Theoretical Results, and Additional Conditions for Theoretical Results, $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$. Define $f$ as the true functional relationship between the response and the unpermuted predictor matrix for all $\omega\in\Omega$,
$\mathbf{Y}_{\omega}=f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)+\bm{\mathscr{U}}_{\omega},$ (45)
and the approximating model form for permutations $m=1,\ldots,M$ and featurizations $r=1,\ldots,\mathscr{R}$.
$\mathbf{Y}_{\omega}=\mathbf{H}_{m,\omega,r}\mathbf{W}^{L}_{m,\omega,r}+\mathbf{U}_{m,\omega,r}$ (46)
{condition}
For every $\eta>0$, there exists a fixed $N$, where $1+\gamma d+q+p\leq N<\infty$, such that as the number of random featurizations $\mathscr{R}\rightarrow\infty$,
$\sup_{\omega\in\Omega}\left\|\frac{1}{\mathscr{R}}\sum_{r=1}^{\mathscr{R}}\mathbf{H}_{1,\omega,r}\mathbf{W}^{L}_{1,\omega,r}-f\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\right)\right\|<\eta,$ (47)
where $\mathbf{W}^{L}_{1,\omega,r}$ is the true coefficient matrix of feature space $r$ for the unpermuted covariate set, $\mathbf{H}_{1,\omega,r}=g\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag},\omega}\;\mathbf{Z}_{\omega}\;\tilde{\mathbf{X}}_{1,\omega}\right]\mathbf{W}_{r}\right)$.
Note that Condition Additional Conditions for Theoretical Results specifically pertains to the $f$ piece of the true functional relationship; no assumption is made of the closeness of a transformed response $\mathbf{H}_{1,\omega,r}\mathbf{W}^{L}_{1,\omega,r}$ and the true values $\mathbf{Y}_{\omega}$ if the predictors themselves are not reliable and the entries of $\bm{\mathscr{U}}_{\omega}$ in Equation 45 are large.
Finite Sample Distribution
For a finite sample, the distribution of the estimated quantity $\hat{\vartheta}_{m}$ is derived as a sum of linear combinations of chi-square random variables. Define the following chi-square random variables
for all $\omega\in\Omega_{\text{obs}}$, $r=1,\ldots,\mathscr{R}$, $k=1,\ldots,K$, $i=1,\ldots,T_{k}d$ and $j<i$. Denote
$\mathbf{H}_{m,\omega,r,k}\left[\mathbf{H}_{m,\omega,r,-k}^{\prime}\mathbf{H}_{m,\omega,r,-k}\right]^{-1}\mathbf{H}_{m,\omega,r,k}^{\prime}=(\phi_{m,\omega,r,k,ij})$ (49)
and $\mathbf{S}_{m,\omega,r}=(s_{m,\omega,r,ij})$, with $i^{\prime}=\lceil i/d\rceil$, $j^{\prime}=\lceil j/d\rceil$, $i^{*}=i\bmod d$, and $j^{*}=j\bmod d$.
Theorem 6.1.
Under the listed conditions, a finite sample containing $T$ observations, $\mathscr{R}$ random generated FNNs, and $\varphi$ realizations in the set $\Omega_{\text{obs}}$, the estimate for the variation parameter $\hat{\vartheta}_{m}$ defined in Equation 13 follows a generalized chi-square distribution.
Proofs
Lemma 6.2.
Under the listed conditions, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$.
The proof of Lemma 6.2 begins with the estimate for one test set with a single individual randomly generated FNN given the variation measure $\mathbf{S}_{m,\omega,r}$. We aggregate these expectations over the sets $k=1,\ldots,K$, generated FNNs $r=1,\ldots,\mathscr{R}$ and the $\varphi\geq 1$ observed datasets in $\Omega_{\text{obs}}$.
Proof 6.3 (Proof of Lemma 6.2).
@@
We examine the matrices $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ under Conditions Additional Conditions for Theoretical Results, Additional Conditions for Theoretical Results, Additional Conditions for Theoretical Results, and Additional Conditions for Theoretical Results. Recall $T_{k}$ is the number of rows in the test set, and $T_{-k}$ the number in the training set. We obtain a fixed and finite $N$ from Condition Additional Conditions for Theoretical Results, and $\text{rank}(\mathbf{H}_{m,\omega,r,-k})=N$ under Equation 40 from Condition Additional Conditions for Theoretical Results. We label the singular values of the matrices $\sigma_{i}(\mathbf{H}_{m,\omega,r,-k})>0$ and $\sigma_{i}(\mathbf{H}_{m,\omega,r,k})\geq 0$ for $i=1,\ldots,N$. For any matrix $\mathbf{P}$, $\left\|\mathbf{P}\right\|_{F}=\sqrt{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|p_{ij}|^{2}}=\sqrt{\sum_{i=1}^{\text{rank}(\mathbf{P})}\sigma_{i}(\mathbf{P})^{2}}$. From Condition Additional Conditions for Theoretical Results, each entry in $\mathbf{H}_{m,\omega,r,-k}$ and $\mathbf{H}_{m,\omega,r,k}$ lies in a bounded interval, $h_{m,\omega,r,-k,ij},h_{m,\omega,r,k,ij}\in[a,b]$ such that $|h_{m,\omega,r,-k,ij}|,|h_{m,\omega,r,k,ij}|\leq G$, and we use this to establish the upper bounds for the maximal singular values shown in Equations 66 and 67. To write the minimum values, we define the average squared matrix entry as in Condition Additional Conditions for Theoretical Results, and note the combination of Conditions Additional Conditions for Theoretical Results, Additional Conditions for Theoretical Results, and Additional Conditions for Theoretical Results yields $0<\nu^{2}\leq 1$, $1\leq\xi^{2}<\infty$, and $1\leq\varrho^{4}<\infty$.
@@
Under Condition Additional Conditions for Theoretical Results, $\mathbf{H}_{m,\omega,r,-k}$ has a finite condition number $\kappa\left(\mathbf{H}_{m,\omega,r,-k}\right)\leq\kappa_{\max}$, and we can use the bounds in Equations 66 and 67 to obtain bounds for the minimum singular values. The rank of $\mathbf{H}_{m,\omega,r,k}$ is not necessarily $N$, as $T_{k}$ can be less than $N$, but it is at least $\min\left\{T_{k},N\right\}$.
@@
We return to the expectation equation given in Equation 51 with the two bounds from Equations 82 and 91.
@@
Thus, $\lim_{T\rightarrow\infty}\mathbb{E}\left[\hat{\vartheta}_{m}\right]=\vartheta_{m}$.
Lemma 6.4.
Under the listed conditions, $\lim_{T\rightarrow\infty}\text{Var}\left(\hat{\vartheta}_{m}\right)=\varphi^{-1}\tau_{\omega}^{2}+\left(\varphi\mathscr{R}\right)^{-1}\tau_{r}^{2}$.
For use in the proof of Lemma 6.4, we first state and prove Remark 6.5, and then proceed like in the proof of Lemma 6.2.
Remark 6.5.
@@
with the last line equal to $(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$. Thus, $\text{tr}^{2}\left(\mathbf{P}\right)\leq(n^{2}-n+1)\;\text{tr}\left(\mathbf{P}^{2}\right)$.
Proof 6.7 (Proof of Lemma 6.4).
We isolate the quantity $\text{tr}\left(\mathbf{R}^{\prime}_{m,\omega,r,k}\mathbf{R}_{m,\omega,r,k}\right)$ like in the proof of Lemma 6.2 and examine the conditional variance $\text{Var}\left(\text{tr}\left[\mathbf{R}_{m,\omega,r,k}^{\prime}\mathbf{R}_{m,\omega,r,k}\right]|\mathbf{S}_{m,\omega,r}\right)$. The unconditional variance follows from the law of total variance and from the definition in Equation 13.